# Low-energy interactions between doubly charmed baryons and Goldstone bosons from lattice QCD

### v1

@@ 2 Lattice setup
The results presented in this paper are based on the gauge configurations generated by the CLQCD Collaboration with $2+1$ dynamical quark flavors using the tadpole improved tree-level Symanzik gauge action and Clover fermion action CLQCD:2023sdb . We use four ensembles with the same lattice spacing $a=0.07746$ fm and two different pion masses $M_{\pi}\simeq 305$ MeV and $M_{\pi}\simeq 207$ MeV. Details of the ensembles are listed in Table 1. Among these ensembles, F32P21/F48P21 and F32P30/F48P30 are two couples that share the same pion mass but have different volumes to obtain more kinematic points in the finite-volume spectra, thereby rendering a more robust determination of the scattering parameters. The action of the valence charm quark is the Fermilab action El-Khadra:1996wdx , which controls discretization errors of $\mathcal{O}(am_{c})^{n}$. The tuning of the parameters in the Fermilab action follows the method applied in Ref. Liu:2009jc .
ID $\beta$ aa(fm) $am_{l}$ $am_{s}$ $M_{\pi}$(MeV) $L^{3}\times T$ $N_{\rm cfgs.}$
F32P30 $6.41$ $0.07746(18)$ 0.2295-0.2295 0.2050-0.2050 $303.9(0.6)$ $32^{3}\times 96$ $750$
F48P30 $6.41$ $0.07746(18)$ 0.2295-0.2295 0.2050-0.2050 $304.9(0.4)$ $48^{3}\times 96$ $359$
F32P21 $6.41$ $0.07746(18)$ 0.2320-0.2320 0.2050-0.2050 $208.1(1.9)$ $32^{3}\times 64$ $459$
F48P21 $6.41$ $0.07746(18)$ 0.2320-0.2320 0.2050-0.2050 $207.4(0.7)$ $48^{3}\times 96$ $222$
Table 1: Details of the four ensembles used in this work, including ensemble identifier (ID), gauge coupling $\beta$, lattice spacing $a$, dimensionless bare quark mass parameters for light and strange quarks ($am_{l}$, $am_{s}$), pion mass $M_{\pi}$, lattice volume $L^{3}\times T$, and number of gauge configurations $N_{\rm cfgs.}$.
Quark propagators are computed using the distillation quark smearing method HadronSpectrum:2009krc . The smearing operator is composed of a small number ($N_{\rm ev}$) of the eigenvectors associated with the $N_{\rm ev}$ lowest eigenvalues of the three-dimensional Laplacian on the HYP-smeared gauge field. The number of eigenvectors $N_{\rm ev}$ is $100$ for all four ensembles. We have verified that increasing $N_{\rm ev}$ to $200$ for the large-volume ensembles (F48P30 and F48P21) yields consistent energy levels. The statistical uncertainty is estimated using the bootstrap method with $4000$ samples.
@@ 3.1 Interpolating operators
where $C=i\gamma_{2}\gamma_{0}$ is the charge conjugation matrix and the positive parity projector is defined by $P_{+}=(1+\gamma_{0})/2$. Only the color indices are shown explicitly, while the Dirac indices are suppressed to adhere to the vector/matrix notation. For a hadron state of definite spatial momentum $\mathbf{p}$, the corresponding interpolator can be generically expressed as
@@
with $\mathcal{O}_{h}(\mathbf{x},t)\equiv\mathcal{O}_{h}(x)$ given by Eq. (1) for mesons and Eq. (2) for baryons. On the lattice, the momentum takes values of $\mathbf{p}=(n_{x},n_{y},n_{z})\times(2\pi/L)$ with $n_{x,y,z}$ being integers.
It is well acknowledged that the two-hadron states in the continuum can be classified by the irreducible representations of the SU(2) group, which are denoted by the total angular momentum $J=0,1/2,1,\cdots$. However, the continuum SU(2) group is reduced to the discrete double covering group $O^{(2)}$ of the cubic group $O$ on the lattice. When parity is taken into account, the $O^{(2)}$ is further promoted to $O^{(2)}_{h}=O^{(2)}\otimes Z_{2}$, with $Z_{2}=\{E={I}^{2},{I}\}$ incorporating the spatial inversion I{I}. Consequently, the construction of two-hadron operators from the one-hadron interpolators in eq. (3) should be carried out according to the irrep $\Gamma=\{A_{1}^{\pm},A_{2}^{\pm},E^{\pm},T_{1}^{\pm},T_{2}^{\pm},G_{1}^{\pm},G_{2}^{\pm},H^{\pm}\}$ of the discrete group $O^{(2)}_{h}$. A variety of approaches, such as the partial-wave Berkowitz:2015eaa ; Wallace:2015pxa ; Prelovsek:2016iyo , projection Gockeler:2012yj ; Prelovsek:2016iyo , and helicity methods Dudek:2012gj ; Prelovsek:2016iyo , have been proposed for constructing lattice operators that describe the scattering of particles with arbitrary spin.
@@
Here, $\alpha$ is the Dirac index of the doubly charmed baryon field. The coefficients $C_{\alpha,\mathbf{p}_{\mathbf{1}},\mathbf{p}_{\mathbf{2}}}$ are chosen to ensure the operators transform under the irrep $G_{1}^{-}$. In Table 2, we list the values of $C_{\alpha,\mathbf{p}_{\mathbf{1}},\mathbf{p}_{\mathbf{2}}}$ for the lowest three momenta $p^{2}=0,1,2$ (in units of $(2\pi/L)^{2}$).
$p=0$ 11 $(0,0,0)$ $(0,0,0)$ 11
$p=1$ 1 (1,0,0)(-1,0,0) $(1,0,0)$ 11
11 $(1,0,0)$ (1,0,0)(-1,0,0) 11
11 (0,1,0)(0,-1,0) $(0,1,0)$ 11
11 $(0,1,0)$ (0,1,0)(0,-1,0) 11
11 (0,0,1)(0,0,-1) $(0,0,1)$ 11
11 $(0,0,1)$ (0,0,1)(0,0,-1) 11
$p=\sqrt{2}$ 11 (1,1,0)(-1,-1,0) $(1,1,0)$ 11
11 $(1,1,0)$ (1,1,0)(-1,-1,0) 11
11 (1,0,1)(-1,0,-1) $(1,0,1)$ 11
11 $(1,0,1)$ (1,0,1)(-1,0,-1) 11
11 (0,1,1)(0,-1,-1) $(0,1,1)$ 11
11 $(0,1,1)$ (0,1,1)(0,-1,-1) 11
11 (1,1,0)(-1,1,0) (1,1,0)(1,-1,0) 11
11 (1,1,0)(1,-1,0) (1,1,0)(-1,1,0) 11
11 (1,0,1)(-1,0,1) (1,0,1)(1,0,-1) 11
11 (1,0,1)(1,0,-1) (1,0,1)(-1,0,1) 11
11 (0,1,1)(0,1,-1) (0,1,1)(0,-1,1) 11
11 (0,1,1)(0,-1,1) (0,1,1)(0,1,-1) 11
@@ 3.3 Finite-volume spectra of the two-hadron systems
where $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$ are given in eq. (5), and the subscripts $i,j$ label different values of momenta $p$. In the case that $p=\{0,1,\sqrt{2}\}(2\pi/L)$, the correlation function is a $3\times 3$ matrix. The quark contraction diagrams for the four channels, $\Omega_{cc}\bar{K}^{(-2,1/2)}$, $\Xi_{cc}K^{(1,1)}$, $\Xi_{cc}K^{(1,0)}$ and $\Xi_{cc}\pi^{(0,3/2)}$, required in our simulation, are shown in figures 12, 13, 14 and 15 of appendix C, respectively.
The finite-volume energy spectra of two-hadron systems can be determined through a systematic analysis of the correlation function matrices $\mathbf{C}$ using the variational technique suggested in ref. Luscher:1990ck . Specifically, the generalized eigenvalue problem (GEVP) Luscher:1990ck ; Blossier:2009kd is solved with a fixed time slice $t_{0}$. In this study, we choose $t_{0}=3$ in lattice units and have verified that the resulting energies are stable against variations of this choice:
@@ 4.1 Lüscher’s formula and effective range expansion
The Lüscher quantization condition relates the finite-volume energy spectra, modified by interactions, to the infinite-volume scattering amplitudes. For a general multi-channel system, the quantization condition in the CM frame reads
$\operatorname{det}[\mathbf{1}+i\mathbf{\Sigma}\cdot\mathbf{T}\cdot(1+i\mathbf{M})]=0\ ,$ (11)
where $\mathbf{T}$ is the infinite-volume T-matrix and $\mathbf{M}$ is a matrix of known functions of the dimensionless scattering momentum $q_{i}=p_{i}L/2\pi$ with $i$ being the channel index. For $S$-wave interaction $\mathbf{M}$ is diagonal. Furthermore, $\mathbf{\Sigma}={\rm diag}\{\cdots,\rho_{i},\cdots\}$ is a diagonal matrix with the phase-space factors $\rho_{i}=2p_{i}/\sqrt{s}$ as elements.
In our single-channel case, the $S$-wave scattering amplitude can be parameterized as
@@
where $\delta_{0}$ denotes the $S$-wave phase shift. Note that we employ a normalization such that the $S$-matrix element is associated with the $T$ matrix by $S=1+2ipT$ and the unitarity relation reads ${\rm Im}\,T=p\,|T|^{2}$. Furthermore, with this normalization, $T$ is identical to the amplitude $f_{0^{+}}$ specified by eq. (2.26) of ref. Liang:2023scp , i.e., $T=f_{0^{+}}$. Consequently, the quantization condition of eq. (11) reduces to the well-known form
@@
where the Lüscher Zeta function is defined by Luscher:1990ux
$\mathcal{Z}_{00}(s;q^{2})=\frac{1}{\sqrt{4\pi}}\sum_{\mathbf{n}\in\mathbb{Z}^{3}}\frac{1}{(|\mathbf{n}|^{2}-q^{2})^{s}}\ .$ (14)
Since $p$ is obtainable from the interacting energy levels via eq. (10), one may readily obtain the phase shift $\delta_{0}(p)$ using eq. (13).
@@
where the threshold parameters $a_{0}$, $r_{0}$ and $v_{n}$ ($n\geq 2$) are referred to as the scattering length, effective range and shape parameters, in order. In the next subsection, we are going to determine the scattering lengths of the four single channels with the help of eqs. 13 and 15.
@@ 4.2 Scattering lengths
where $E^{\rm cont.}_{h}(p)$ ($h\in\{\phi,B_{cc}\}$) are the energies calculated from the continuum dispersion relation $E^{\rm cont.}_{h}(p)=\sqrt{m_{h}^{2}+p^{2}}$, while $E^{\rm latt.}_{h}(p)$ corresponds to the single-particle energies computed on the lattice. Such a shift is implemented to alleviate the lattice artifacts arising from discretization effects following Padmanath:2022cvl ; Prelovsek:2020eiw ; Piemonte:2019cbi ; Xing:2022ijm . The other point concerns the selection of the energy levels. That is, only the energy levels below the $B_{c}D$ thresholds are kept in our scattering analysis, as already explained at the end of section 3.3. Our final results of $p\cot\delta_{0}$, derived from energy levels with the dispersion relation correction (DRC) implemented via eq. (16), are shown as red squares and blue dots in figure 2.
@@
Here, the indices $i,j$ run over all energy levels included in the fit for a given scattering channel and pion mass. The quantity $p^{2}(E_{i})$ is the squared momentum determined from the $i$-th lattice energy level $E_{i}$ by numerically solving eq. (10), $p_{\rm ERE}^{2}(i,a_{0},r_{0},\dots)$ is the solution of the the Lüscher quantization condition (13) with ERE parametrization  (15), and $C_{\rm cov}$ is the covariance matrix of the squared momenta.
By truncating the ERE at $\mathcal{O}(p^{2})$ (denoted as leading order (LO) hereafter), we obtain the scattering length $a_{0}$ and effective range $r_{0}$. The resulting values of $a_{0}$ and $r_{0}$ are collected in table 5. The labels “Raw” and “DRC” in the table denote the $p\cot\delta_{0}$ datasets derived from energy levels without and with DRC, respectively. It can be found that the impact of DRC on scattering lengths $a_{0}$ is negligible, since it is predominantly determined by the lowest-lying two-particle energy levels nearest to the non-interacting threshold. Those lowest energy levels have zero momentum $p=\left|\mathbf{p}_{\mathbf{1},\mathbf{2}}\right|=0$, and the DRC term $c^{2}p^{2}$ in the dispersion relation vanishes. In contrast, the effective range $r_{0}$ exhibits pronounced sensitivity to DRC across nearly all scattering channels. For instance, in the $\Xi_{cc}\pi^{(0,\frac{3}{2})}$ system at $M_{\pi}\sim 300$ MeV, the DRC shifts $r_{0}$ from 0.23(11)-0.23(11) fm to 0.03(19)-0.03(19) fm with $\chi^{2}/{\rm dof}$ improved from $4.29$ to $1.54$. We adopt the values from fits to the DRC-applied data as our final results for both threshold parameters, although $a_{0}$ is robust against the DRC. Based on the fitted values, our LO ERE predictions of $p\cot\delta_{0}$ are presented as solid green lines with bands in figure 2.
@@
In the above, the $S$-wave scattering lengths for the four distinct channels have been determined at two unphysical pion masses, $\sim$ 300 and 210 MeV. To extrapolate them to the physical pion mass $M_{\pi}^{\mathrm{phy}}=135$ MeV, we employ a rank-one polynomial in $M_{\pi}^{2}$ as follows:
@@
The central values of the extrapolated results are obtained by performing the linear fit (18) to the scattering lengths from the LO ERE analysis with DRC applied. The first uncertainty quoted is the statistical error (sta) propagated from the $1$-$\sigma$ uncertainties of the fit parameters. We estimate two dominant sources of systematic uncertainty, both originating from the choice of input values for the extrapolation. The first systematic uncertainty (sys1) is associated with the DRC, which we quantify as the difference between the central extrapolated value and that obtained using the Raw dataset. The second systematic uncertainty (sys2) stems from the truncation of the ERE, estimated as the difference when using the results from the NLO fits as input for the extrapolation. Note that for the $\Xi_{cc}K^{(1,0)}$ channel, the systematic uncertainty from the ERE truncation is not quoted, as an NLO fit was not feasible due to the limited number of data points.

### v2

@@ 2 Lattice setup
The results presented in this paper are based on the gauge configurations generated by the CLQCD Collaboration with $2+1$ dynamical quark flavors using the tadpole improved tree-level Symanzik gauge action and Clover fermion action CLQCD:2023sdb . We use four ensembles with the same lattice spacing $a=0.07746$ fm and two different pion masses $M_{\pi}\simeq 305$ MeV and $M_{\pi}\simeq 207$ MeV. Details of the ensembles are listed in Table 1. Among these ensembles, F32P21/F48P21 and F32P30/F48P30 are two couples that share the same pion mass but have different volumes to obtain more kinematic points in the finite-volume spectra, thereby rendering a more robust determination of the scattering parameters. The action of the valence charm quark is the Fermilab action El-Khadra:1996wdx , which controls discretization errors of $\mathcal{O}(am_{c})^{n}$. The tuning of the parameters in the Fermilab action follows the method applied in Ref. Liu:2009jc . Specifically, the charm quark mass parameter $m_{c}$ and the anisotropy parameter $\nu$ in the action are tuned by requiring the mass of $J/\psi$ to reproduce the physical value and the dispersion relation to match the continuum form.
ID $\beta$ aa(fm) $am_{l}$ $am_{s}$ $M_{\pi}$(MeV) $L^{3}\times T$ $M_{\pi}L$ $N_{\rm cfgs.}$
F32P30 $6.41$ $0.07746(18)$ 0.2295-0.2295 0.2050-0.2050 $303.9(0.6)$ $32^{3}\times 96$ $3.818(8)$ $750$
F48P30 $6.41$ $0.07746(18)$ 0.2295-0.2295 0.2050-0.2050 $304.9(0.4)$ $48^{3}\times 96$ $5.746(8)$ $359$
F32P21 $6.41$ $0.07746(18)$ 0.2320-0.2320 0.2050-0.2050 $208.1(1.9)$ $32^{3}\times 64$ $2.614(24)$ $459$
F48P21 $6.41$ $0.07746(18)$ 0.2320-0.2320 0.2050-0.2050 $207.4(0.7)$ $48^{3}\times 96$ $3.908(13)$ $222$
Table 1: Details of the four ensembles used in this work, including ensemble identifier (ID), gauge coupling $\beta$, lattice spacing $a$, dimensionless bare quark mass parameters for light and strange quarks ($am_{l}$, $am_{s}$), pion mass $M_{\pi}$, lattice volume $L^{3}\times T$, dimensionless spatial size $M_{\pi}L$, and number of gauge configurations $N_{\rm cfgs.}$.
Quark propagators are computed using the distillation quark smearing method HadronSpectrum:2009krc . The smearing operator is composed of a small number ($N_{\rm ev}$) of the eigenvectors associated with the $N_{\rm ev}$ lowest eigenvalues of the three-dimensional Laplacian on the HYP-smeared gauge field. The number of eigenvectors $N_{\rm ev}$ is $100$ for all four ensembles. We have verified that increasing $N_{\rm ev}$ to $200$ for the large-volume ensembles (F48P30 and F48P21) yields consistent energy levels. The statistical uncertainty is estimated using the bootstrap method with $4000$ samples. We have verified that this sample size is sufficient for the convergence of both the mean values and their statistical uncertainties.
@@ 3.1 Interpolating operators
The interpolating operators for the Goldstone bosons ($\pi^{+}$, $K^{+}$ and $K^{0}$) and the ground-state $(1/2)^{+}$ DCBs ($\Xi_{cc}^{++,+}$, $\Omega_{cc}^{+}$) employed in this work are standard, which can be found in, e.g., refs. Gattringer:2010zz ; Liu:2009jc . For easy reference and completeness, we provide their explicit forms in Appendix B. In what follows, we focus on the construction of the two-hadron operators specific to our simulation.
For a hadron state of definite spatial momentum $\mathbf{p}$, the corresponding interpolator can be generically expressed as
@@
with $\mathcal{O}_{h}(\mathbf{x},t)\equiv\mathcal{O}_{h}(x)$ given by eq. (19) for mesons and eq. (20) for baryons. On the lattice, the momentum takes values of $\mathbf{p}=(n_{x},n_{y},n_{z})\times(2\pi/L)$ with $n_{x,y,z}$ being integers. It is well acknowledged that the two-hadron states in the continuum can be classified by the irreducible representations of the SU(2) group, which are denoted by the total angular momentum $J=0,1/2,1,\cdots$. However, the continuum SU(2) group is reduced to the discrete double covering group $O^{(2)}$ of the cubic group $O$ on the lattice. When parity is taken into account, the $O^{(2)}$ is further promoted to $O^{(2)}_{h}=O^{(2)}\otimes Z_{2}$, with $Z_{2}=\{E={I}^{2},{I}\}$ incorporating the spatial inversion I{I}. Consequently, the construction of two-hadron operators from the one-hadron interpolators in eq. (1) should be carried out according to the irrep $\Gamma=\{A_{1}^{\pm},A_{2}^{\pm},E^{\pm},T_{1}^{\pm},T_{2}^{\pm},G_{1}^{\pm},G_{2}^{\pm},H^{\pm}\}$ of the discrete group $O^{(2)}_{h}$. A variety of approaches, such as the partial-wave Berkowitz:2015eaa ; Wallace:2015pxa ; Prelovsek:2016iyo ; Yan:2025jlq , projection Gockeler:2012yj ; Prelovsek:2016iyo ; Yan:2025jlq , and helicity methods Dudek:2012gj ; Prelovsek:2016iyo , have been proposed for constructing lattice operators that describe the scattering of particles with arbitrary spin.
@@
Here, $\alpha$ is the Dirac index of the doubly charmed baryon field. The coefficients $C_{\alpha,\mathbf{p}_{\mathbf{1}},\mathbf{p}_{\mathbf{2}}}$ are chosen to ensure the operators transform under the irrep $G_{1}^{-}$. In Table 6 of Appendix B , we list the values of $C_{\alpha,\mathbf{p}_{\mathbf{1}},\mathbf{p}_{\mathbf{2}}}$ for the lowest three momenta $p^{2}=0,1,2$ (in units of $(2\pi/L)^{2}$).
@@ 3.3 Finite-volume spectra of the two-hadron systems
where $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$ are given in eq. (3), and the subscripts $i,j$ label different values of momenta $p$. For the set of momenta $p=\{0,1,\sqrt{2}\}(2\pi/L)$, the correlation function is a $3\times 3$ matrix. The quark contraction diagrams for the four channels, $\Omega_{cc}\bar{K}^{(-2,1/2)}$, $\Xi_{cc}K^{(1,1)}$, $\Xi_{cc}K^{(1,0)}$ and $\Xi_{cc}\pi^{(0,3/2)}$, required in our simulation, are shown in figures 12, 13, 14 and 15 of appendix C, respectively.
The finite-volume energy spectra of two-hadron systems can be determined through a systematic analysis of the correlation function matrices $\mathbf{C}$ using the variational technique suggested in ref. Luscher:1990ck . Specifically, the generalized eigenvalue problem (GEVP) Michael:1985ne ; Luscher:1990ck ; Blossier:2009kd is solved with a fixed time slice $t_{0}$. In this study, we choose $t_{0}=3$ in lattice units and have verified that the resulting energies are stable against variations of this choice:
@@ 4.1 Lüscher’s formula and effective range expansion
The Lüscher quantization condition relates the finite-volume energy spectra, modified by interactions, to the infinite-volume scattering amplitudes. In our single-channel case, the $S$-wave scattering amplitude can be parameterized as
@@
where $\delta_{0}$ denotes the $S$-wave phase shift. Note that we employ a normalization such that the $S$-matrix element is associated with the $T$ matrix by $S=1+2ipT$ and the unitarity relation reads ${\rm Im}\,T=p\,|T|^{2}$. Furthermore, with this normalization, $T$ is identical to the amplitude $f_{0^{+}}$ specified by eq. (2.26) of ref. Liang:2023scp , i.e., $T=f_{0^{+}}$. Consequently, the Lüscher quantization condition has the well-known form
@@
where $\mathcal{Z}_{00}(s;q^{2})$ is the Lüscher Zeta function defined in ref. Luscher:1990ux . Since $p$ is obtainable from the interacting energy levels via eq. (8), one may readily obtain the phase shift $\delta_{0}(p)$ using eq. (10).
@@
where the threshold parameters $a_{0}$, $r_{0}$ and $v_{n}$ ($n\geq 2$) are referred to as the scattering length, effective range and shape parameters, in order. In the next subsection, we are going to determine the scattering lengths of the four single channels with the help of eqs. 10 and 11.
@@ 4.2 Scattering lengths
where $E^{\rm cont.}_{h}(p)$ ($h\in\{\phi,B_{cc}\}$) are the energies calculated from the continuum dispersion relation $E^{\rm cont.}_{h}(p)=\sqrt{m_{h}^{2}+p^{2}}$, while $E^{\rm latt.}_{h}(p)$ corresponds to the single-particle energies computed on the lattice. Such a shift is implemented to alleviate the lattice artifacts arising from discretization effects following Padmanath:2022cvl ; Prelovsek:2020eiw ; Piemonte:2019cbi ; Xing:2022ijm . Specifically, since the Lüscher formalism assumes continuum kinematics, the deviation of the lattice dispersion relation from the continuum form (where $c^{2}\neq 1$) can introduce systematic errors. This correction aligns the kinematic behavior of the lattice data with the continuum expectation, thereby ensuring consistency with the quantization condition. The other point concerns the selection of the energy levels. That is, only the energy levels below the $B_{c}D$ thresholds are kept in our scattering analysis, as already explained at the end of section 3.3. Our final results of $p\cot\delta_{0}$, derived from energy levels with the dispersion relation correction (DRC) implemented via eq. (12), are shown as red squares and blue dots in figure 2.
@@
Here, the indices $i,j$ run over all energy levels included in the fit for a given scattering channel and pion mass. The quantity $p^{2}(E_{i})$ is the squared momentum determined from the $i$-th lattice energy level $E_{i}$ by numerically solving eq. (8), $p_{\rm ERE}^{2}(i,a_{0},r_{0},\dots)$ is the solution of the Lüscher quantization condition (10) with ERE parametrization  (11), and $C_{\rm cov}$ is the covariance matrix of the squared momenta.
By truncating the ERE at $\mathcal{O}(p^{2})$ (denoted as leading order (LO) hereafter), we obtain the scattering length $a_{0}$ and effective range $r_{0}$. The resulting values of $a_{0}$ and $r_{0}$ are collected in table 4. The labels “Raw” and “DRC” in the table denote the $p\cot\delta_{0}$ datasets derived from energy levels without and with DRC, respectively. It can be found that the Raw and DRC-applied results are consistent with each other within 1-$\sigma$ uncertainties, indicating the impact of DRC on $a_{0}$ and $r_{0}$ is mild. In particular for $a_{0}$, it is predominantly determined by the lowest-lying two-particle energy levels nearest to the non-interacting threshold, which have zero momentum $p=\left|\mathbf{p}_{\mathbf{1},\mathbf{2}}\right|=0$ with vanishing DRC term $c^{2}p^{2}$ in the dispersion relation. Nevertheless, we adopt the values from fits to the DRC-applied data as our final results for both threshold parameters. Based on the fitted values, our LO ERE predictions of $p\cot\delta_{0}$ are presented as solid green lines with bands in figure 2.
@@
In the above, the $S$-wave scattering lengths for the four distinct channels have been determined at two unphysical pion masses, $\sim$ 300 and 210 MeV. To extrapolate them to the physical pion mass $M_{\pi}^{\mathrm{phy}}=135$ MeV, we employ a rank-one polynomial in $M_{\pi}^{2}$ inspired by BChPT as follows:
@@
The central values of the extrapolated results are obtained by performing the linear fit (14) to the scattering lengths from the LO ERE analysis with DRC applied. The first uncertainty quoted is the statistical error (sta) propagated from the $1$-$\sigma$ uncertainties of the fit parameters. We estimate two dominant sources of systematic uncertainty, both originating from the choice of input values for the extrapolation. The first systematic uncertainty (sys1) is associated with the DRC, which we quantify as the difference between the central extrapolated value and that obtained using the Raw dataset. The second systematic uncertainty (sys2) stems from the truncation of the ERE, estimated as the difference when using the results from the NLO fits as input for the extrapolation. It is worth noting that this error budget represents a rough estimation based on current data. To achieve a robust control over the systematic uncertainties, calculations at more pion masses and lattice spacings are mandatory. Note that for the $\Xi_{cc}K^{(1,0)}$ channel, the systematic uncertainty from the ERE truncation is not quoted, as an NLO fit was not feasible due to the limited number of data points.